If the coefficient of static friction is 0.23 and the coefficient of kinematics friction is 0.18 what force is required to start a 8.6 kg object moving ?
1
Expert's answer
2019-04-24T09:45:35-0400

Applying the Newton’s Second Law of Motion, we get (at the moment when the object begins to move the acceleration is zero):


"\sum F_x = ma_x = 0,""F_{appl} - F_{s.fr.} = 0,""F_{appl} = F_{s.fr.} = \mu_s N = \mu_s mg,"

here, "F_{appl}" is the applied force required to start a 8.6 kg object to move, "F_{s.fr.}" is the force of static friction, "\mu_s = 0.23" is the coefficient of static friction, "N" is the normal force, "m = 8.6 kg" is the mass of the object and "g" is the acceleration due to gravity.

Then, we get:


"F_{appl} = \mu_s mg = 0.23 \cdot 8.6 kg \cdot 9.8 \dfrac{m}{s^2} = 19.4 N."

Answer:

"F_{appl} = 19.4 N."